20x+5(x+1)= 25(x+1)-20
I came up with 20=20x
x=1
This is not one of my choices
no solution
any real number
x=0
x=5

Answer by stanbon(75887)

20x+5(x+1)= 25(x+1)-20
20x+5x+5=25x+25-20
25x+5=25x+5
x=x
OR
0=0
This tells you the equation is
an identity. It is true for any value
of x you would care to put in it.
It is equivalent to saying x=x
cheers,
Stan H.